what is the answer to this 5n+34=-2(1-7n)
Answer by checkley79(3341)

5n+34=-2(1-7n)
5n+34=-2+14n
5n-14n=-2-34
-9n-36
n=-36/-9
n=4 ans.
Proof:
5*4+34=-2(1-7*4)
20+34=-2(1-28)
54=-2(-27)
54=54